# Contrastive Instruction Tuning

### v1

@@ Abstract
Instruction tuning has been used as a promising approach to improve the performance of large language models (LLMs) on unseen tasks. However, current LLMs exhibit limited robustness to unseen instructions, generating inconsistent outputs when the same instruction is phrased with slightly varied forms or language styles. This behavior indicates LLMs’ lack of robustness to textual variations and generalizability to unseen instructions, potentially leading to trustworthiness issues. Accordingly, we propose Contrastive Instruction Tuning (Coin), which maximizes the similarity between the hidden representations of semantically equivalent instruction-instance pairs while minimizing the similarity between semantically different ones. To facilitate this approach, we augment the existing FLAN collection by paraphrasing task instructions. Experiments on the PromptBench benchmark show that Coin consistently improves LLMs’ robustness to unseen instructions with variations across character, word, sentence, and semantic levels by an average of $+2.5\%$ in accuracy.
@@ 1 Introduction
In this paper, we propose Contrastive Instruction Tuning (Coin), an instruction tuning method that leverages contrastive learning to align the hidden representations of instruction-instance pairs that are semantically equivalent but textually different and to differentiate those that are semantically distinct. Given the same input and output instance, we pair each instruction with its perturbed versions as positive samples. Observed that the hidden representations of data from different tasks already have low cosine similarity with each other Liu et al. (2023a), we use the same instruction paired with different instance input and output as hard negative samples (refer to Section 3.2 for more details). Intuitively, by recognizing that the same instruction with different formulations can have the same meaning, the model can generate more consistent answers given different instructions of the same intent and become more robust to variations in language expressions. At the same time, negative samples encourage the model to understand that an instruction can lead to different outcomes in different contexts, facilitating the model to generate more desired outputs at the instance level.
@@ 3.1 Overview
As explored by many previous studies, instruction-tuning with large-scale datasets mainly focuses on aligning the desired output for a given instruction-instance pair from various tasks (Sanh et al., 2022; Longpre et al., 2023; Wei et al., 2022). However, current LLMs exhibit a lack of robustness to the same instruction with varying forms (Sun et al., 2023; Zhu et al., 2023; Liang et al., 2023). To mitigate this limitation, our method Coin further leverages contrastive learning to maximize the similarity between hidden representations of semantically equivalent instruction-instance pairs. This approach enhances models’ robustness and consistency to variations in instruction expressions.
@@ 3.2 Contrastive Data Selection
For negative samples, we observe that the contrastive loss converges quickly when using instruction-input pairs of different tasks (i.e., normal negatives), leading to minor improvement in robustness. This observation is consistent with the findings in prior studies (Liu et al., 2023a): LLMs can distinguish between instructions of different tasks such that their hidden representations already have low cosine similarity. To collect hard negatives, we draw inspiration from near-OOD samples, which are data that come from the same task but with different classes Winkens et al. (2020); Fort et al. (2021); Liu et al. (2023a). Prior studies found that it is more difficult for models to detect near-OOD samples than samples from other tasks (far-OOD). This finding indicates that the hidden representations of near-OOD samples may not be distinguishable enough and thus can provide informative supervision signals for contrastive learning. Accordingly, we choose such a sample $(I_{i}^{-},x_{i}^{-},y_{i}^{-})$ that shares the same instruction as the original instance ($I_{i}^{-}=I_{i}$) but is paired with different input ($x_{i}^{-}\neq x_{i}$) and output ($y_{i}^{-}\neq y_{i}$) as a negative sample.
@@ 4.1 Training Datasets
In this work, we conduct experiments on a widely used instruction tuning dataset: the FLAN Collection Wei et al. (2022). FLAN Collection Wei et al. (2022) is a large-scale data collection that encompasses a wide range of tasks, including natural language inference, common sense reasoning, sentiment analysis, paraphrase identification, etc. This data collection is created by transforming 62 publicly available text datasets into instructional formats. $10$ unique instruction templates are manually composed for each dataset. In this work, we choose 25 datasets with deterministic answers from the collection. Iterating through the training split of each dataset in a round-robin manner and for each entry in the dataset, we create a positive sample and select one entry from the remaining entries from the dataset as a negative sample following the selection strategy described in Section 3.2. Refer to Appendix A for more details of the processed dataset.
@@ 4.3 Evaluation Setting
Instruction Variations. Regarding instructions, we select six clean instructions predefined for each task. Then, we create perturbed version of each instruction. Following PromptBench, we use DeepWordBug Gao et al. (2018) to introduce character-level perturbations for certain words, and use TextFooler Jin et al. (2020) to replace words with contextually similar words. At the sentence level, we implement the CheckList Ribeiro et al. (2020) and append randomly generated sequences, which all have a length of $10$ and consist of alphabets and digits, at the end of an instruction to distract LLMs. For the semantic-level perturbation, PromptBench defines $10$ prompts for each task following the linguistic behavior of six languages: Chinese, French, Arabic, Spanish, Japanese, and Korean. To keep the number of instructions the same as other types of perturbation, we randomly select one instruction from each language defined for each task, which are all different from the clean prompts.
@@ 4.4 Results
While continual instruction tuning on tasks similar to the testing data increases the model’s accuracy in general, the performance gap between different groups can still be as large as $6.1\%$ observed between the clean group and the group with word-level perturbation. This shows that the continually instruction-tuned model still lacks robustness to unseen instructions with variations across different levels.
Compared to the base model and continual instruction tuning, Coin significantly reduces performance variance and consistently improves accuracy for instructions within and across different groups. As it can be seen from Figure 3, Coin achieves improvements in accuracy for all types of perturbation, up to $4.4\%$ for word-level perturbations where the continually instruction-tuned model exhibits its largest drop in performance. The largest performance gap is reduced to $3.6\%$. The consistent improvement across all types of perturbations demonstrates the generalizability of Coin at enhancing models’ robustness against variations in instructions at different levels. Coin also decreases the performance variance on instructions from the five groups by $1.6\%$, $1.9\%$, $2.1\%$, $2.5\%$, and $1.2\%$. This also shows that Coin can effectively help the model become less sensitive to specific instructions for each task and more consistent in its performance.
@@ 5.2 Impact on Different Tasks
Table 1: Models’ average accuracy of each task category. Coin has consistent improvement across all tasks with more evident improvement on duplicate sentence detection and grammar correctness tasks.
@@
On the other hand, Coin gains modest enhancement in accuracy on sentiment analysis and NLI tasks by +1.4% and +1.7% compared to the continual instruction-tuned model. For the sentiment analysis task, the continually instruction-tuned model has already achieved an accuracy of 89.0%. Obtaining further improvements can be challenging given that the model is already capable at distinguishing between sentences with different sentiments. Regarding NLI, the task requires a comprehensive understanding of the relationship between two sentences, which can depend on the model’s knowledge of various domains or reasoning ability to infer implicit meanings that are not directly stated. The complex relation between two sentences may not be explicitly captured by the hidden representations, meaning that Coin may not explicitly further enhance the model’s reasoning ability. However, Coin still obtains an improvement of +1.4% and +1.7% on the two tasks, demonstrating Coin’s effectiveness at enhancing the model’s ability to discern the nuanced emotion cues or inferential relation that underlies the overall semantic meaning of the instruction-input pairs.
@@ Appendix B Detailed Experiment Results
DeepWordBug 61.8 $\pm$ 4.6 47.2 $\pm$ 6.4 47.2 $\pm$ 6.4 49.3 $\pm$ 4.5 27.4 $\pm$ 24.1 42.7 $\pm$ 6.9 15.6 $\pm$ 10.9 55.5 $\pm$ 5.6 66.7 $\pm$ 15.6 49.3 $\pm$ 3.5 46.3 $\pm$ 18.0
TextFooler 61.7 $\pm$ 2.0 49.6 $\pm$ 3.8 49.6 $\pm$ 3.8 49.2 $\pm$ 4.7 43.3 $\pm$ 21.8 24.8 $\pm$ 17.4 14.7 $\pm$ 8.1 57.5 $\pm$ 4.9 46.4 $\pm$ 25.8 53.1 $\pm$ 2.7 45.0 $\pm$ 18.7
CheckList 64.8 $\pm$ 1.8 51.2 $\pm$ 3.6 51.2 $\pm$ 3.6 52.9 $\pm$ 2.2 15.3 $\pm$ 10.7 50.2 $\pm$ 3.2 22.6 $\pm$ 6.8 61.5 $\pm$ 3.3 82.3 $\pm$ 4.1 52.1 $\pm$ 2.0 50.4 $\pm$ 19.0
Continual Instruction-Tuned Clean 63.5 $\pm$ 8.6 68.7 $\pm$ 2.4 67.3 $\pm$ 2.7 66.3 $\pm$ 2.7 62.8 $\pm$ 13.0 62.9 $\pm$ 4.2 71.2 $\pm$ 7.6 82.0 $\pm$ 1.9 90.1 $\pm$ 2.4 57.5 $\pm$ 3.8 69.2 $\pm$ 11.1
DeepWordBug 64.9 $\pm$ 3.1 64.9 $\pm$ 2.1 64.1 $\pm$ 2.3 63.4 $\pm$ 1.9 62.1 $\pm$ 11.9 54.7 $\pm$ 3.6 61.9 $\pm$ 11.8 75.7 $\pm$ 4.8 90.5 $\pm$ 2.0 54.0 $\pm$ 5.1 65.6 $\pm$ 11.7
TextFooler 58.9 $\pm$ 12.6 64.8 $\pm$ 4.1 65.4 $\pm$ 3.8 64.3 $\pm$ 3.5 56.4 $\pm$ 10.5 46.8 $\pm$ 6.7 62.5 $\pm$ 8.2 73.8 $\pm$ 3.5 84.2 $\pm$ 12.6 54.0 $\pm$ 2.1 63.1 $\pm$ 12.6
CheckList 58.6 $\pm$ 15.2 66.4 $\pm$ 1.9 65.3 $\pm$ 1.4 65.1 $\pm$ 3.7 55.9 $\pm$ 16.8 53.2 $\pm$ 8.6 66.6 $\pm$ 8.1 80.3 $\pm$ 3.0 90.4 $\pm$ 1.2 55.9 $\pm$ 4.3 65.8 $\pm$ 13.9
@@
The detailed results on different datasets is shown in Table 3. For the results of models trained with different contrastive loss weighting, refer to Table 4.

### v2

@@ Abstract
Instruction tuning has been used as a promising approach to improve the performance of large language models (LLMs) on unseen tasks. However, current LLMs exhibit limited robustness to unseen instructions, generating inconsistent outputs when the same instruction is phrased with slightly varied forms or language styles. This behavior indicates LLMs’ lack of robustness to textual variations and generalizability to unseen instructions, potentially leading to trustworthiness issues. Accordingly, we propose Contrastive Instruction Tuning (Coin), which maximizes the similarity between the hidden representations of semantically equivalent instruction-instance pairs while minimizing the similarity between semantically different ones. To facilitate this approach, we augment the existing FLAN collection by paraphrasing task instructions. Experiments on the PromptBench benchmark show that Coin consistently improves LLMs’ robustness to unseen instructions with variations across character, word, sentence, and semantic levels by an average of $+2.5\%$ in accuracy. [^1] Code is available at https://github.com/luka-group/CoIN.
@@ 1 Introduction
In this paper, we propose Contrastive Instruction Tuning (Coin), an instruction tuning method that leverages contrastive learning to align the hidden representations of instruction-instance pairs that are semantically equivalent but textually different and to differentiate those that are semantically distinct. Given the same input and output instance, we pair each instruction with its perturbed versions as positive samples. Observed that the hidden representations of data from different tasks already have low cosine similarity with each other Liu et al. (2023a), we use the same instruction paired with different instance input and output as hard negative samples (refer to Section 3.2 for more details). Intuitively, by recognizing that the same instruction with different formulations can have the same meaning, the model can generate more consistent answers given different instructions of the same intent and become more robust to variations in language expressions. At the same time, negative samples encourage the model to understand that an instruction can lead to different outcomes in different contexts, facilitating the model to distinguish inputs with different user intents.
@@ 3.1 Overview
As explored by many previous studies, instruction-tuning with large-scale datasets mainly focuses on aligning the desired output for a given instruction-instance pair from various tasks (Sanh et al., 2022; Longpre et al., 2023; Wei et al., 2022). However, current LLMs exhibit a lack of robustness when facing the same instruction expressed in different forms (Sun et al., 2023; Zhu et al., 2023; Liang et al., 2023), causing LLMs to be unreliable when being deployed in the real world. To mitigate this limitation, our method Coin further leverages contrastive learning to maximize the similarity between hidden representations of semantically equivalent instruction-instance pairs. This approach enhances models’ robustness and consistency to variations in instruction expressions.
@@ 3.2 Contrastive Data Selection
For negative samples, we observe that the contrastive loss converges quickly when using instruction-input pairs of different tasks (i.e., normal negatives), leading to minor improvement in robustness. This observation is consistent with the findings in prior studies (Liu et al., 2023a): LLMs can distinguish between instructions of different tasks such that their hidden representations already have low cosine similarity. To collect hard negatives, we draw inspiration from near-OOD samples, which are data that come from the same task but with different classes Winkens et al. (2020); Fort et al. (2021); Liu et al. (2023a). Prior studies found that it is more difficult for models to detect near-OOD samples than samples from other tasks (far-OOD). This finding indicates that the hidden representations of near-OOD samples may not be distinguishable enough and thus can provide informative supervision signals for contrastive learning. Accordingly, we choose such a sample $(I_{i}^{-},x_{i}^{-},y_{i}^{-})$ that shares the same instruction as the original instance ($I_{i}^{-}=I_{i}$) but is paired with different input ($x_{i}^{-}\neq x_{i}$) and output ($y_{i}^{-}\neq y_{i}$) as a negative sample. For example, if $y_{i}$ is "yes", then $y_{i}^{-}$ can be "no", ensuring the fundamental intent of the instruction-instance pair is different from the original one. Based on this approach, Coin encourages the model to align semantically equivalent instructions with different phrasings while contrasting inputs with different user intents.
@@ 4.1 Training Datasets
In this work, we conduct experiments on a widely used instruction tuning dataset: the FLAN Collection Wei et al. (2022). FLAN Collection Wei et al. (2022) is a large-scale data collection that encompasses a wide range of tasks, including natural language inference, common sense reasoning, sentiment analysis, paraphrase identification, etc. This data collection is created by transforming 62 publicly available text datasets into instructional formats. $10$ unique instruction templates are manually composed for each dataset. In this work, we choose 25 datasets with deterministic answers from the collection. To ensure each dataset has an equal chance of being sampled into the training set of Coin, we iterate through the training split of each dataset with a round-robin approach. For each entry, we create a positive sample by randomly selecting a predefined instruction template not used by the entry to paraphrase the instruction. Only paraphrasing is used for creating training data while various types of perturbations are included for evaluation (refer to Section 4.3). Avoiding assumptions about specific types of noise in instructions is crucial due to the high uncertainty LLMs face in real-world deployment. Hence, a robustness training method that can generalize to other types of perturbations is more desirable. We then select one entry from the remaining dataset as a negative sample, following the strategy in Section 3.2. Refer to Appendix A for more details of the processed dataset.
@@ 4.3 Evaluation Setting
Instruction Variations. Regarding instructions, we select six clean instructions predefined for each task. Then, we create perturbed versions of each instruction. Following PromptBench, we use DeepWordBug Gao et al. (2018) to introduce character-level perturbations for certain words, and use TextFooler Jin et al. (2020) to replace words with contextually similar words. At the sentence level, we implement the CheckList Ribeiro et al. (2020) and append randomly generated sequences, which all have a length of $10$ and consist of alphabets and digits, at the end of an instruction to distract LLMs. For the semantic-level perturbation, PromptBench defines $10$ instructions that paraphrase the original instruction for each task while following the linguistic behavior of six languages: Chinese, French, Arabic, Spanish, Japanese, and Korean. To keep the number of instructions the same as other types of perturbation, we randomly select one instruction from each language defined for each task, which are all different from the clean instructions. We also ensure that instructions used for evaluation differ from all instructions in the training dataset and thus are unseen by the model, preventing data contamination.
@@ 4.4 Results
Compared to the base model, the continually instruction-tuned model shows increases in accuracy, which is expected as the model is exposed to more data and trained with more steps. Nevertheless, the performance gap between different groups can still be as large as $6.1\%$ observed between the clean group and the group with word-level perturbation. This shows that the continually instruction-tuned model still lacks robustness to unseen instructions with variations across different levels.
Compared to continual instruction tuning, Coin further reduces performance variance and consistently improves accuracy for instructions within and across different groups without introducing any new data and training steps. As it can be seen from Figure 3, Coin achieves improvements in accuracy for all types of perturbation, up to $4.4\%$ for word-level perturbations where the continually instruction-tuned model exhibits its largest drop in performance. The largest performance gap is reduced to $3.6\%$. The consistent improvement across all types of perturbations demonstrates the generalizability of Coin at enhancing models’ robustness against variations in instructions at different levels. Coin also decreases the performance variance on instructions from the five groups by $1.6\%$, $1.9\%$, $2.1\%$, $2.5\%$, and $1.2\%$. This also shows that Coin can effectively help the model become less sensitive to specific instructions for each task and more consistent in its performance. For more detailed results, refer to Table 2.
@@ 5.2 Impact on Different Tasks
Table 1: Models’ average accuracy and standard deviation of each task category. Coin has consistent improvement across all tasks with more evident improvement on duplicate sentence detection and grammar correctness tasks.
@@
On the other hand, Coin gains modest enhancement in accuracy on sentiment analysis and NLI tasks by +1.4% and +1.7% compared to the continual instruction-tuned model. For the sentiment analysis task, the continually instruction-tuned model has already achieved an accuracy of 89.0%. Obtaining further improvements can be challenging given that the model is already capable at distinguishing between sentences with different sentiments. Regarding NLI, the task requires a comprehensive understanding of the relationship between two sentences, which can depend on the model’s knowledge of various domains or reasoning ability to infer implicit meanings that are not directly stated. The complex relation between two sentences may not be explicitly captured by the hidden representations, meaning that Coin may not explicitly further enhance the model’s reasoning ability. However, Coin still obtains an improvement of +1.4% and +1.7% on the two tasks, demonstrating Coin’s effectiveness at enhancing the model’s ability to discern the nuanced inferential relation that underlies the overall semantic meaning of the instruction-input pairs.
Character 61.8 $\pm$ 4.6 47.2 $\pm$ 6.4 47.2 $\pm$ 6.4 49.3 $\pm$ 4.5 27.4 $\pm$ 24.1 42.7 $\pm$ 6.9 15.6 $\pm$ 10.9 55.5 $\pm$ 5.6 66.7 $\pm$ 15.6 49.3 $\pm$ 3.5 46.3 $\pm$ 18.0
Word 61.7 $\pm$ 2.0 49.6 $\pm$ 3.8 49.6 $\pm$ 3.8 49.2 $\pm$ 4.7 43.3 $\pm$ 21.8 24.8 $\pm$ 17.4 14.7 $\pm$ 8.1 57.5 $\pm$ 4.9 46.4 $\pm$ 25.8 53.1 $\pm$ 2.7 45.0 $\pm$ 18.7
Sentence 64.8 $\pm$ 1.8 51.2 $\pm$ 3.6 51.2 $\pm$ 3.6 52.9 $\pm$ 2.2 15.3 $\pm$ 10.7 50.2 $\pm$ 3.2 22.6 $\pm$ 6.8 61.5 $\pm$ 3.3 82.3 $\pm$ 4.1 52.1 $\pm$ 2.0 50.4 $\pm$ 19.0
Continual Instruction Tuning Clean 63.5 $\pm$ 8.6 68.7 $\pm$ 2.4 67.3 $\pm$ 2.7 66.3 $\pm$ 2.7 62.8 $\pm$ 13.0 62.9 $\pm$ 4.2 71.2 $\pm$ 7.6 82.0 $\pm$ 1.9 90.1 $\pm$ 2.4 57.5 $\pm$ 3.8 69.2 $\pm$ 11.1
Character 64.9 $\pm$ 3.1 64.9 $\pm$ 2.1 64.1 $\pm$ 2.3 63.4 $\pm$ 1.9 62.1 $\pm$ 11.9 54.7 $\pm$ 3.6 61.9 $\pm$ 11.8 75.7 $\pm$ 4.8 90.5 $\pm$ 2.0 54.0 $\pm$ 5.1 65.6 $\pm$ 11.7
Word 58.9 $\pm$ 12.6 64.8 $\pm$ 4.1 65.4 $\pm$ 3.8 64.3 $\pm$ 3.5 56.4 $\pm$ 10.5 46.8 $\pm$ 6.7 62.5 $\pm$ 8.2 73.8 $\pm$ 3.5 84.2 $\pm$ 12.6 54.0 $\pm$ 2.1 63.1 $\pm$ 12.6
Sentence 58.6 $\pm$ 15.2 66.4 $\pm$ 1.9 65.3 $\pm$ 1.4 65.1 $\pm$ 3.7 55.9 $\pm$ 16.8 53.2 $\pm$ 8.6 66.6 $\pm$ 8.1 80.3 $\pm$ 3.0 90.4 $\pm$ 1.2 55.9 $\pm$ 4.3 65.8 $\pm$ 13.9
Character 66.9 $\pm$ 3.0 68.2 $\pm$ 2.0 67.5 $\pm$ 1.3 66.6 $\pm$ 4.0 72.4 $\pm$ 2.5 58.7 $\pm$ 4.2 64.7 $\pm$ 8.0 78.5 $\pm$ 3.1 91.1 $\pm$ 2.1 58.9 $\pm$ 2.6 69.4 $\pm$ 9.8
Word 66.5 $\pm$ 4.5 67.4 $\pm$ 1.7 67.7 $\pm$ 3.0 66.1 $\pm$ 2.3 71.9 $\pm$ 5.4 49.9 $\pm$ 7.5 63.9 $\pm$ 6.0 75.6 $\pm$ 3.5 85.6 $\pm$ 11.6 60.1 $\pm$ 3.8 67.5 $\pm$ 10.5
Sentence 68.4 $\pm$ 7.2 67.7 $\pm$ 3.5 68.2 $\pm$ 2.6 66.3 $\pm$ 3.6 63.3 $\pm$ 9.6 55.4 $\pm$ 9.5 66.8 $\pm$ 6.1 79.8 $\pm$ 3.5 92.3 $\pm$ 0.6 59.6 $\pm$ 2.8 68.8 $\pm$ 11.4
@@ 6 Conclusion
Acknowledgement
We appreciate the reviewers for their insightful comments and suggestions. Tianyi Yan was supported by the Center for Undergraduate Research in Viterbi Engineering (CURVE) Fellowship. Fei Wang was supported by the Amazon ML Fellowship. James Y. Huang was supported by a gift fund from the Amazon Center on Secure & Trusted ML. Muhao Chen was supported by the NSF Grant IIS 2105329, the NSF Grant ITE 2333736, the DARPA AIE Grant HR0011-24-9-0370, and an Amazon Research Award.
@@ Appendix B Detailed Experiment Results
For the results of models trained with different contrastive loss weighting, refer to Table 4.